OpenStudy (anonymous):

Solve. 5x+2/2=2x-1 x =

OpenStudy (anonymous):

it is x= -2/3

OpenStudy (anonymous):

I am so sorry...

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

I cant put that...

OpenStudy (anonymous):

why

OpenStudy (anonymous):

I can only put in 3 different numbers

OpenStudy (anonymous):

then do it in decimal form

OpenStudy (anonymous):

but -2/3 is the answer though...

OpenStudy (anonymous):

that is true @Electric_Blue

OpenStudy (anonymous):

5x+2/2=2x-1 We simplify the equation to the form, which is simple to understand 5x+2/2=2x-1 We move all terms containing x to the left and all other terms to the right. +5x-2x=-1-1 We simplify left and right side of the equation. +3x=-2 We divide both sides of the equation by 3 to get x. x=-0.66666666666667

OpenStudy (anonymous):

isn't like -0.667

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

it will only let me put in -0.

OpenStudy (anonymous):

I am right though :-)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

-2/3 is the answer....

OpenStudy (anonymous):

I don't know how you will put it in.. it can written in either decimal or fraction

OpenStudy (anonymous):

what kind of test or exam is that? that it only lets you put in a certain number of numbers?

OpenStudy (anonymous):

exactly!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Give me medal and fan

OpenStudy (anonymous):

I helped u

OpenStudy (anonymous):

@Electric_Blue I gave u medal too!

OpenStudy (anonymous):

you are very supportive to the answer

OpenStudy (anonymous):

thanks for that

OpenStudy (anonymous):

Yea you too

OpenStudy (anonymous):

So what do I do

OpenStudy (anonymous):

do you have a teacher to ask for questions?

OpenStudy (anonymous):

Nope Im on k12

OpenStudy (anonymous):

and that is?

OpenStudy (anonymous):

Home Schooling

OpenStudy (anonymous):

oh but its not online? like im in flvs

OpenStudy (anonymous):

No it's online

OpenStudy (anonymous):

so its a different online school that you dont have teachers to help you with, am i right?

OpenStudy (anonymous):

No then can help but there not online

OpenStudy (anonymous):

I like getting ahead so I can get done before school is over

OpenStudy (anonymous):

they*

OpenStudy (anonymous):

ah ok well that stinks, is there anyway you can call or text them

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

well i dont know how to help you, because if i was in that situation i would contact my teacher right away

OpenStudy (anonymous):

Okay well thank you for the help you did do! :)

OpenStudy (anonymous):

your welcome hope you get help soon👋

OpenStudy (radar):

Maybe it is copied down wrong???? May be it is supposed to be